# Exploring LLMs for Scientific Information Extraction using the SciEx Framework

### v1

@@ Introduction
Current approaches to applying LLMs to scientific information extraction use fine-tuning (dunn2022structured), prompt engineering (polak2024extracting; da2024automated), or in-context learning. However, despite promising results, building stable and generalizable pipelines remains difficult due to prompt sensitivity, terminological and numerical inconsistency, and long-document dependencies. Scientific papers frequently exceed model context windows and distribute related evidence across multiple modalities, challenging current architectures to integrate this information coherently.
@@
The modular and extensible design of SciEx allows flexible replacement, extension, and integration of off-the-shelf components. As a prompt-driven framework, SciEx maintains adaptability by offering interfaces for researchers to incorporate domain-specific prompt designs when beneficial. In addition, SciEx supports automatic prompt optimization via tools such as DSPy (khattab2023dspy), which automate prompt engineering to enhance extraction performance.
@@
PDF Extractor. The first step of SciEx involves a PDF Extractor parsing scientific publications in PDF format to extract both textual and visual content for downstream information extraction. In our implementation, we employ Docling [^1] https://docling-project.github.io/docling/, an open-source document conversion toolkit that performs fine-grained layout analysis and structural recognition (livathinos2025docling).
@@
We primarily evaluate SciEx using two large language models: Gemini-2.5-Flash (comanici2025gemini) and GPT-4o (hurst2024gpt), applied for both textual and visual information extraction in our experiments. Importantly, the framework is model-agnostic; other LLMs, VLMs or MLLMs can be substituted for specific modality extraction tasks. For retrieval, we retrieve the top-5 most relevant chunks from the contextualized database each round.
@@
Large language models (LLMs) (brown2020language; ouyang2022training) have advanced text understanding, reasoning, and generation, inspiring generative information extraction (zhang-etal-2025-survey), where tasks like NER and RE are reformulated as sequence generation. Instruction-tuned and few-shot LLMs (wei2021finetuned; kojima2022large) enable flexible schema adaptation without retraining, while domain-specific variants (lee2020biobert; chithrananda2020chemberta; gu2021domain) enhance understanding of specialized corpora. RAG (lewis2020retrieval) further improves factuality through evidence grounding.
LLMs for Scientific Literature. Early work emphasized text-centric understanding through classification (vajjala2025text), summarization (zhang2024comprehensive), citation recommendation, and entity–relation extraction (huo2019semi). Domain applications include materials science (olivetti2020data), joint entity–relation extraction (dagdelen2024structured), and medical IE using weak supervision (das2025weakly). More recent studies demonstrate that LLMs achieve strong zero- and few-shot extraction without fine-tuning (ghosh-etal-2024-toward; da2024automated; woo2024evaluation; bhayana2024use), while ChatExtract (polak2024extracting) enables conversational, iterative extraction. Since much scientific knowledge resides in figures and tables, multi-modal understanding has gained attention. Prior work on figure analysis, chart captioning (hsu-etal-2021-scicap-generating; tang-etal-2023-vistext; yang2024scicap_plus; kantharaj-etal-2022-chart), and table reasoning (zhou-etal-2025-m2) often relied on task-specific models with costly annotations. Recent multi-modal LLMs (MLLMs) such as mPLUG-DocOwl (hu2024mplug), mPLUG-1.5 (hu2024mplug1_5), and UReader (ye-etal-2023-ureader) unify text–image reasoning in shared semantic spaces, though they remain limited to local contexts. OmniParser (wan2024omniparser) integrates text spotting, key information extraction, and table recognition, yet long-document reasoning and structured output generation remain challenging.
LLM-powered Tools for Scientific Information Extraction. A growing ecosystem of LLM-based tools supports interactive literature exploration. ChatPDF [^2] https://www.chatpdf.com/ and ChatDoc [^3] https://www.chatdoc.com/ enable conversational querying of papers. Elicit [^4] https://elicit.com/ and SciSpace [^5] https://scispace.com/extract-data/ automate summarization, data extraction, and multi-paper comparison. ScholarPhi (head2021augmenting) and Qlarify (fok2023qlarify) enhance interpretability via linked definitions and recursive exploration. LangExtract [^6] https://langextract.com/ employs grounded extraction with visualization, while SciDaSynth (wang2025scidasynth) combines automated extraction with human validation for reliability.
@@ Acknowledgement
Older or poorly formatted PDFs present significant challenges because of their degraded visual quality and inconsistent layouts. Graphs in these documents are frequently scanned or rasterized at low resolutions, often with truncated axes, unreadable scales or ambiguous values (e.g., Figure 5, image is from (harper1963influence)) which hinders the LLM’s ability to accurately extract numerical data.
@@
Figure 6: Example table from (chaudhuri1970removal)
Figure 7: Example table from (abbaszadegan2008removal)
@@
Figure 8: Clustered data points. Example figure from (chang1985uv)
Figure 9: Single legend. Example figure from (fedorenko2020survival)
Figure 10: Large figure with many subfigures. Example figure from (van2020aerosol)
@@
Figure 11: Cross referencing. Example from (yang2012relationship)

### v2

@@ Introduction
Current approaches to applying LLMs to scientific information extraction use fine-tuning (Dunn et al. 2022), prompt engineering (Polak and Morgan 2024; da Silva et al. 2024), or in-context learning. However, despite promising results, building stable and generalizable pipelines remains difficult due to prompt sensitivity, terminological and numerical inconsistency, and long-document dependencies. Scientific papers frequently exceed model context windows and distribute related evidence across multiple modalities, challenging current architectures to integrate this information coherently.
@@
The modular and extensible design of SciEx allows flexible replacement, extension, and integration of off-the-shelf components. As a prompt-driven framework, SciEx maintains adaptability by offering interfaces for researchers to incorporate domain-specific prompt designs when beneficial. In addition, SciEx supports automatic prompt optimization via tools such as DSPy (Khattab et al. 2023), which automate prompt engineering to enhance extraction performance.
@@
PDF Extractor. The first step of SciEx involves a PDF Extractor parsing scientific publications in PDF format to extract both textual and visual content for downstream information extraction. In our implementation, we employ Docling [^1] https://docling-project.github.io/docling/, an open-source document conversion toolkit that performs fine-grained layout analysis and structural recognition (Livathinos et al. 2025).
@@
We primarily evaluate SciEx using two large language models: Gemini-2.5-Flash (Comanici et al. 2025) and GPT-4o (Hurst et al. 2024), applied for both textual and visual information extraction in our experiments. Importantly, the framework is model-agnostic; other LLMs, VLMs or MLLMs can be substituted for specific modality extraction tasks. For retrieval, we retrieve the top-5 most relevant chunks from the contextualized database each round.
@@
Large language models (LLMs) (Brown et al. 2020; Ouyang et al. 2022) have advanced text understanding, reasoning, and generation, inspiring generative information extraction (Zhang et al. 2025), where tasks like NER and RE are reformulated as sequence generation. Instruction-tuned and few-shot LLMs (Wei et al. 2021; Kojima et al. 2022) enable flexible schema adaptation without retraining, while domain-specific variants (Lee et al. 2020; Chithrananda et al. 2020; Gu et al. 2021) enhance understanding of specialized corpora. RAG (Lewis et al. 2020) further improves factuality through evidence grounding.
LLMs for Scientific Literature. Early work emphasized text-centric understanding through classification (Vajjala and Shimangaud 2025), summarization (Zhang et al. 2024), citation recommendation, and entity–relation extraction (Huo et al. 2019). Domain applications include materials science (Olivetti et al. 2020), joint entity–relation extraction (Dagdelen et al. 2024), and medical IE using weak supervision (Das et al. 2025). More recent studies demonstrate that LLMs achieve strong zero- and few-shot extraction without fine-tuning (Ghosh et al. 2024; da Silva et al. 2024; Woo et al. 2024; Bhayana et al. 2024), while ChatExtract (Polak and Morgan 2024) enables conversational, iterative extraction. Since much scientific knowledge resides in figures and tables, multi-modal understanding has gained attention. Prior work on figure analysis, chart captioning (Hsu et al. 2021; Tang et al. 2023; Yang et al. 2024; Kantharaj et al. 2022), and table reasoning (Zhou et al. 2025) often relied on task-specific models with costly annotations. Recent multi-modal LLMs (MLLMs) such as mPLUG-DocOwl (Hu et al. 2024a), mPLUG-1.5 (Hu et al. 2024b), and UReader (Ye et al. 2023) unify text–image reasoning in shared semantic spaces, though they remain limited to local contexts. OmniParser (Wan et al. 2024) integrates text spotting, key information extraction, and table recognition, yet long-document reasoning and structured output generation remain challenging.
LLM-powered Tools for Scientific Information Extraction. A growing ecosystem of LLM-based tools supports interactive literature exploration. ChatPDF [^2] https://www.chatpdf.com/ and ChatDoc [^3] https://www.chatdoc.com/ enable conversational querying of papers. Elicit [^4] https://elicit.com/ and SciSpace [^5] https://scispace.com/extract-data/ automate summarization, data extraction, and multi-paper comparison. ScholarPhi (Head et al. 2021) and Qlarify (Fok et al. 2023) enhance interpretability via linked definitions and recursive exploration. LangExtract [^6] https://langextract.com/ employs grounded extraction with visualization, while SciDaSynth (Wang et al. 2025) combines automated extraction with human validation for reliability.
@@ Acknowledgement
Older or poorly formatted PDFs present significant challenges because of their degraded visual quality and inconsistent layouts. Graphs in these documents are frequently scanned or rasterized at low resolutions, often with truncated axes, unreadable scales or ambiguous values (e.g., Figure 5, image is from (Harper 1963)) which hinders the LLM’s ability to accurately extract numerical data.
@@
Figure 6: Example table from (Chaudhuri and Engelbrecht 1970)
Figure 7: Example table from (Abbaszadegan et al. 2008)
@@
Figure 8: Clustered data points. Example figure from (Chang et al. 1985)
Figure 9: Single legend. Example figure from (Fedorenko et al. 2020)
Figure 10: Large figure with many subfigures. Example figure from (Van Doremalen et al. 2020)
@@
Figure 11: Cross referencing. Example from (Yang et al. 2012)